# Anonymization Prompt Learning for Facial Privacy-Preserving Text-to-Image Generation

### v1

@@ 4.1 Setup
Dataset. We curate dataset $S_{ID}$ of image-prompt pairs for training our Anonymization Prompt. We first collect names of identities that encompass different genders, races, and occupations. With these names as prompts, training images $x^{i}$s are generated with Stable Diffusion v1.5 ldm . We next use ChatGPT chatgpt to generate $c_{2}$ prompts by filling in names and requesting descriptions of the individual’s demographic characteristics and prominent facial attributes. While these responses potentially include inaccurate descriptions, we find this experimental collection procedure to be quite effective in our experiments. We additionally collect another dataset of identities $S_{test}$ for evaluating anonymization on unseen identities. Samples of our collected dataset can be found in the Appendix.
Training. The Anonymization Prompt (AP) is trained on the base model of Stable Diffusion v1.5 ldm , an advanced text-to-image diffusion model. Hyperparameters are set as $m=10$ and $\alpha=1$. During training, we set the learning rate at $31e-$ 3, and train for $20,000$ steps with batch sizes of 1. We keep all model parameters frozen and only train the learnable prompt.
Evaluating Anonymization. We measure the identity accuracy (ID-ACC) of generated images by calculating the cosine similarity between its Arcface arcface face recognition features and that of the ground truth identity. Lower identity accuracies imply better anonymization. Besides evaluating on SD-v1.5, transferability of AP is tested on two other text-to-image models, Stable Diffusion v1.4 ldm and RealisticVision v5.1 realisticvision . The former is an earlier version of Stable Diffusion, while the latter is a fine-tuned version of Stable Diffusion that emphasizes realistic, detailed photo generation.
Evaluating Quality. We assess generation quality using captions from MS COCO 2014 evaluation set  coco , following common practices in generative modeling research ldm ; imagen ; dalle3 . We report CLIP clip and FID cleanfid scores to respectively evaluate semantic and quality preservation. We evaluate the preservation of facial attribute with a facial attribute recognizer based on FaRL farl .
@@ 4.2 Anonymization across Models
Table 1: The effect of the Anonymization Prompt (AP) on identity accuracies of three text-to-image models. All models generate highly accurate identities when prompted with names of public figures. Employing AP significantly reduces accuracies. We calculate average / maximum identity accuracies for images generated from the same identity, and aggregate these metrics across all identities.
@@
SD-v1.5 ldm 0.31566 / 0.37958 0.30254 / 0.36691
SD-v1.5 ldm + APL 0.03924 / 0.10861 0.03970 / 0.10773
SD-v1.4 ldm 0.29911 / 0.36713 0.28837 / 0.35720
SD-v1.4 ldm + APL 0.03893 / 0.10632 0.04647 / 0.11368
RV-v5.1 realisticvision 0.32205 / 0.37122 0.32071 / 0.36492
RV-v5.1 realisticvision + APL 0.09215 / 0.14391 0.09033 / 0.14496
Figure 3: Images generated by three different models when employing the same Anonymization Prompts. Models are prompted to generate Albert Einstein and Taylor Swift, which respectively belong to the training set and the testing set. The visual effect demonstrate successful anonymization for any identities across different text-to-image models, consistent with our quantitative results.
Following the training of Anonymization Prompt, we prepend the learned prefix to three different text-to-image diffusion models to evaluate its anonymization performances. We prompt the models to generate images of public figures and calculate identity accuracies (ID-ACC). In this experiment, we use the prompt template "A close-up portrait of [Name]" to ensure that the models generate clear depictions of faces. For each identity, we generate a group of images and calculate both average and maximum ID-ACCs within the group. Subsequently, we aggregate these scores across all identities to present overall average and maximum identity accuracies. Results are presented in table 1.
Experiments show that original text-to-image models perform well on generating facial identities, reaching an average ID-ACC of around 0.3. However, with Anonymization Prompts, ID-ACCs drop to as low as 0.03, marking almost a tenfold decrease from the original version. Notably, the anonymization performances on the testing set are on par with those on the training set, indicating an effectiveness to all identities, beyond those available during training. Visually, images generated by all models are significantly less recognizable, as demonstrated in  fig. 3, consistent with our quantitative results. We thereby confirm the validity of Anonymization Prompt on three different models, despite only being trained on SD-v1.5, showcasing a remarkable transferability of our method.
@@ 4.3 Quality Preservation
We strive to ensure our anonymization is natural and specific to identifiable facial images. When requested to generate specific individuals, models equipping Anonymization Prompts should return facial images with false identities, but share similar facial attributes with the requested identity, as outlined in section 3.4. In addition, this modification should not interfere with generation qualities of unrelated images, i.e., concatenating the AP to prompts of non-facial images and non-identity-specific facial images should not significantly change the outputs.
@@
In this section, we comprehensively evaluate the ability of our method to meet these requirements. On top of the aforementioned identity accuracies, we use common image generation evaluators, including FID and CLIP scores to evaluate semantic and quality preservation. For each method, we generate $30k$ images with prompts from COCO-30k coco . We evaluate FID between images generated by modified models and the original Stable Diffusion ldm , and evaluate CLIP score by averaging the distances between all generated images and their corresponding text prompts. For attribute evaluation, we evaluate generated facial images in section 4.2 with Attribute ACC, which is the average cosine similarity between the facial attribute features farl of the generated faces and the ground truths.
@@
We compare our results with ESD erase , Concept Ablation (CA) ablating and Forget-me-not (FMN) forgetmenot , which are recent concept editing methods that also target diffusion models. We run their open training code on SD-v1.5 ldm but substitute their editing concepts with identities in our training set.
Table 2: Quantitative results of semantic, quality and attribute preservation. We report average ID-ACC on the testing sets and other quality measures specified above. Best results are in bold and second bests are underlined.
ID-ACC (↓) Attribute ACC (↑) FID (↓) CLIP (↑) Transferrable
SD-v1.5 ldm 0.3025 0.6936 - 0.3078 -
ESD erase - - 9.4256 0.2530 No
CA ablating 0.0122 0.5389 4.0382 0.2903 No
FMN forgetmenot 0.0792 0.6154 5.8753 0.2783 No
Ours (APL) 0.0397 0.6022 2.7137 0.3059 Yes
@@
Results in table 2 demonstrate that our method best satisfies all requirements for a successful anonymization. APL achieves the highest image quality and text fidelity, respectively quantified by FID and CLIP scores, surpassing all compared methods. In generating facial images, it not only yields low identity accuracies, but also preserves facial attribute information of the requested identities.
In contrast, all other methods fall short under at least one metric. ESD erase fails to generate any facial images after training since ESD erase aligns its target concepts to empty strings, which completely removes targets concepts. Similarly, Concept Ablation ablating selects an anchor concept, such as "a person", as optimization targets. While it generates facial images with satisfying identity accuracies, the images share little facial attribute similarity with the requested identity, as indicated by a markedly lower Attribute ACC. As mentioned in section 3.4, we consider these anonymizations to be excessive, as they produce completely irrelevant faces or no faces at all. We provide visualizations of these anonymizations in Appendix.
@@
APL sets appropriate optimization targets that specifies basic demographic and facial attribute information for all training identities, and therefore preserves necessary attributes while removing identity information. Its regularization loss $L_{Reg}$ effectively preserves generation quality of unrelated images, as we show in the ablation study. Finally, its prompt learning setting not only ensures minimal quality degradation compared to the original model, but also makes it the only method that can be trained once and then transferred to other models.
@@ 4.4 Anonymizing new identities
Personalization methods, such as Dreambooth dreambooth , Textual Inversion textualinversion and HyperDreambooth hyperdreambooth , enable diffusion models to acquire new concepts, including highly accurate facial identities hyperdreambooth . This capability significantly expands deepfake threats, affecting individuals beyond those in the model’s training dataset. Meanwhile, these methods also undermine the effectiveness of traditional concept editing methods erase ; ablating , as studies show that concepts removed by them can be re-introduced to the model with the same personalization procedure imma ; ringabell .
Our experiments demonstrate that the Anonymization Prompt remains effective amid these challenges. Firstly, it effectively anonymizes identities learned with Dreambooth, even though the identities was never seen during training. Secondly, because APL does not modify model parameters, re-learning anonymized identity is not effective. After any attempts of re-learning, users can simply add the AP back on and still generate anonymized images. These capabilities are not preserved by previously compared methods, which unanimously modify model parameters. Detailed experiments can be found in the Appendix.
@@ 4.5 Ablation Study
We study the ablations of several key designs in Anonymization Prompt Learning, including training iterations and regularization. We additionally investigate different choices of prompt length, weight parameter $\alpha$ and training dataset sizes in Appendix.
@@
Figure 4: Effects of training iterations. (a) An identity generated by RealisticVision realisticvision that employs Anonymization Prompts of different training iterations. Different anonymization levels of the same identity is observed. (b) Identity accuracies computed across the test set. Users of AP may choose to use AP at different training iterations according to their own policies of content moderation.
@@
Varying training iterations. We monitor the identity accuracies during the entire training process, and observe that the accuracies for all models progressively decrease as training progresses in fig. 4(a). This effect is visualized with one identity in fig. 4(b). Both visual effects and quantitative measures indicate a steadily increasing depth of anonymization. In light of these findings, text-to-image service providers may choose to employ Anonymization Prompts at various stages of training to enforce different levels of anonymization in alignment with their content moderation policies.
@@
Figure 5: Effect of our regularization loss $L_{Reg}$. When training the AP solely with facial images, the model links AP to the concept of faces, leading to unexpected depictions of people in every generation. Regularization removes the issue by inactivating AP for non-ID prompts.
@@
Regularization. We ablate the regularization loss $L_{Reg}$, which trains the Anonymization Prompt to be ineffective when concatenated to non-identity-specific prompts. We demonstrate in fig. 5 that without regularization, models generate unexpected depictions of people, which is a significant violation of semantic alignment. Since APL is a prompt learning method, training without the regularization of non-face images causes the model to link the prompt to the concept of faces, resulting in above situations. Therefore, training the AP to ensure its ineffectiveness on non-identity-specific prompts is essential for our method to be useful in practice.
5 Conclusions
We propose Anonymization Prompt Learning (APL), which produces a plug-and-play Anonymization Prompt (AP) that can be inserted to any text-to-image diffusion models and enforce the generation of anonymized faces. We show the ability of AP to anonymize various identities on different models while preserving the basic generation performances of the original diffusion model. Put together, AP has the potential to be incorporated by text-to-image service platforms to prevent the generation of identity-specific images and address deepfake generations. We hope the development of AP contributes to the long-term battle against malicious use of AI-generated content.
Appendix
Appendix A Additional Experimental Results
A.1 Anonymizing New Identities
We present the details of the experiments described in  section 4.4 of our paper. These experiments concern the ability of Anonymization Prompts (AP) to take effects on identities learned through personalization methods, such as Dreambooth [27]. We show that AP is resilient to the introduction of new identities, even those learned after the training process of AP.
Figure 6: We evaluate the anonymization performance of our Anonymization Prompt (AP) on identites learned after training AP. (a) Identity accuracies of images generated with and without AP at different Dreambooth [27] iterations. AP consistently lowers identity accuracies on newly-learned identities. (b) An example of anonymizing a new identity [28], which is linked to a token [V] by Dreambooth [27]. AP significantly reduces the identity accuracy of the generated image.
We introduce new identities to Stable Diffusion v1.5 [26] with Dreambooth [27] and evaluate the anonymization performance of AP on these identities. Specifically, we sample 50 identities from the VGGFace-2 [6] dataset, collecting high-resolution photos for each identity for personalization. We obtain a personalized Stable Diffusion for each of those identities with Dreambooth [27], so that resulting models are able to accurately generate their corresponding identities with a pre-determined token. Models are fine-tuned for 800 iterations, which is sufficient for the base model to learn accurate identities. During training, images of the corresponding identities are generated every 100 iteration, both with and without applying AP, to measure their identity accuracies. The results across all identities are then aggregated to form the curves depicted in fig. 6(a).
Owing to the effectiveness of Dreambooth [27], both curves exhibit a clear increasing trend. However, in cases where identities are generated with AP, the identity accuracies consistently remain lower compared to those generated without AP across all iterations. With fewer iterations, AP effectively reduces the identity accuracies to nearly half of the base model. At 800 iteratieons, AP achieves an average identity accuracy of 0.43, significantly lower than the baseline of 0.53, demonstrating effective anonymizations on these new identities when Dreambooth [27] is trained until convergence.
We visualize the effect of AP on one learned identity in fig. 6(b). At the same training iteration of Dreambooth [27], the identity generated without AP closely resembles the ground truth identity, whereas the one generated with AP is much less recognizable, showcasing successful anonymization.
These figures demonstrate both quantitatively and qualitatively that AP effectively anonymizes identities learned by the base model after AP’s training phase. Consequently, modifications to the model parameters to either introduce new identities or restore previously erased ones do not affect anonymization performance. This finding significantly extends the applicability of AP to encompass all possible facial identities, thereby offering protection against deepfake threats and privacy violations not only for public figures but also for the general population.
A.2 Additional Visualizations
Figure 7: Anonymization performance of different optimization targets. Empty string targets of ESD [12] lead to random scenery generations and category label targets of Concept Ablation [19] generate every identity as a random person. Our prompts curated with abundant attribute information strikes a balance between erasure of semantics and attaining original attributes.
We provide visualizations of different optimization targets in accordance to quantitative results in  table 2. As underscored in  section 3.4, we train the Anonymization Prompt to generate individuals of different identities but with similar facial attributes. This is achieved by optimizing the outputs of diffusion models to align with our target prompts, which describes the demographic and facial attribute information of corresponding identities. We demonstrate that these optimization targets are essential for successful anonymizations.
We compare our settings of optimization targets with empty strings and category labels, which are respectively employed by concept editing methods ESD [12] and Concept Ablation [19]. In two identity-specific generation cases in  fig. 7, we show that all three optimization targets achieve anonymization. However, our anonymization yields the best visual effect, preserving key facial attributes of the requested identities. In contrast, the compared methods tend to cause excessive anonymization. To erase the concept of a person, ESD [12] trains its diffusion model to predict noise that align with the textual feature of an empty string. Such empty string targets result in low-quality, random scenery depictions for prompts of specific individuals. Similarly, Concept Ablation [19] uses category labels as optimization targets, such as "a person", which lead to completely random person generation, consistent with the low Attribute ACC of Concept Ablation [19] in table 2. These results of anonymization may appear unnatural for users, potentially leading to poorer user experiences.
A.3 Additional Ablation Studies
Figure 8: We evaluate the identity accuracies (ID-ACC) of Stable Diffusion v1.5, enhanced with Anonymization Prompts (AP) trained on differing sizes of datasets. Zero denotes evaluations without AP. Evaluation is performed on the test set $S_{test}$ distinct from training identities. We document both the mean and maximum ID-ACC for a group of images per identity, then aggregate these metrics across all identities. We observe a consistent decline in both mean and maximum ID-ACC as the number of identities in the training set increases.
Dataset Size. Anonymization Prompt enables the anonymization of any identities, including those unseen during training and those learned through fine-tuning. We show that this generalization ability has its roots in the number of training identities. Results are presented in  fig. 8. With more training identities, we observe a consistent decrease in both average and maximum Identity Accuracies. This suggests that incorporating more identities enables AP to identify more precise directions from $c_{1}$ to $c_{2}$ in the latent space, where $c_{1}$ represents the prompts specifiying identities and $c_{2}$ are prompts encapsulating the demographic information and facial attributes of these identities. An accurate estimation of this direction guides the model to anonymize more identities successfully. We expect the identity accuracies to further drop on unseen diffusion models if larger datasets with more diverse attributes are available.
Prompt length. The length of Anonymization Prompt is the number of learnable token that constitutes our prompt prefix. We study the optimal length of the prompt by evaluating Identity Accuracies with AP of different lengths. As shown in  table 3, we find that $m=10$ strikes a balance between effective anonymization and computational efficiency. Shorter prompts are less effective at anonymization. Longer prompts only offer marginal improvements, while requiring more computational and storage resources.
@@ A.3 Additional Ablation Studies
Table 3: Identity Accuracy of Anonymization Prompt with different prompt length $m$. We report mean / maximum ID-ACC in a group of image for the same individual, and then aggregate across all identities.
@@
Table 4: Identity Accuracy of Anonymization Prompt trained with different weight parameter $\alpha$. We report mean / maximum ID-ACC in a group of image for the same individual, and then aggregate across all identities.
@@
A.4 Samples from collected dataset
We show some samples from our collected dataset $S_{ID}$ in table 5. We reiterate that the attribute information is automatically generated by ChatGPT [1], and they may not accurately reflect the actual characteristics of these individuals.

### v2

@@ 4.1 Setup
Dataset. We curate dataset $S_{ID}=\{(x^{i},c_{1}^{i},c_{2}^{i})\}_{i=1}^{N}$ for training our Anonymization Prompt. We first collect names of identities that encompass different genders, races, and occupations. With these names as prompts, training images $x^{i}$s are generated with Stable Diffusion v1.5 [1]. We use ChatGPT [27] to generate $c_{2}$ prompts by filling in names and requesting descriptions of the individual’s demographic characteristics and prominent facial attributes. While these responses potentially include inaccurate descriptions, we find this experimental collection procedure to be quite effective in our experiments. We additionally collect another dataset of identities $S_{test}$ for evaluating anonymization on unseen identities. Both $S_{ID}$ and $S_{test}$ contain 100 non-overlapping identities of public figures. Samples of these datasets can be found in the Appendix.
Training. The Anonymization Prompt is trained on the base model of Stable Diffusion v1.5 [1] (SD-v1.5), an advanced text-to-image diffusion model. Hyperparameters are set as $m=10$ and $\alpha=1$. During training, we set the learning rate at $31e-$ 3, and train for $20,000$ steps with batch sizes of 1. We keep all model parameters frozen and only train the learnable prompt.
Evaluating Anonymization. We measure the identity accuracy (ID-ACC) of generated images by calculating the cosine similarity between its Arcface [28] face recognition features and that of the ground truth identity. Lower identity accuracies imply better anonymization. Besides evaluating on SD-v1.5, we demonstrate the transferability of our method with three other text-to-image models, namely SD-v1.4 [1], RealisticVision v5.1 [10], and Stable Diffusion XL (SDXL) [14]. Note that SDXL [14] utilizes a text encoder with different sizes of text embeddings, which bars a direct transfer of our soft prompt. We resolve this issue with a zero-shot prompt transfer method proposed in [29], mapping our prompt to a different embedding space, so that our method remains effective with different model architectures.
Evaluating Quality. We assess the generation quality of both identity-specific and non-identity-specific images. For the former, we calculate FID [30] and facial attribute similarities between faces generated by our model and the original model. For the latter, we generate images with captions from the MS COCO 2014 evaluation set [31], following common practices in generative modeling research [1, 3, 2]. We report CLIP [32] and FID [30] scores to respectively evaluate semantic and quality preservation. Given that our method modifies an existing generative model, we calculate FID with reference images generated by the original model in all cases.
@@ 4.2 Anonymization across Models
Table 1: The effect of the Anonymization Prompt on identity accuracies of three text-to-image models. All models generate highly accurate identities when prompted with names of public figures. Employing the prompt significantly reduces accuracies. We calculate mean / maximum identity accuracies for images generated from the same identity, and aggregate these metrics across all identities.
@@
SD-v1.5 [1] 0.31566 / 0.37958 0.30254 / 0.36691
SD-v1.5 [1] + APL 0.03924 / 0.10861 0.03970 / 0.10773
SD-v1.4 [1] 0.29911 / 0.36713 0.28837 / 0.35720
SD-v1.4 [1] + APL 0.03893 / 0.10632 0.04647 / 0.11368
RV-v5.1 [10] 0.32205 / 0.37122 0.32071 / 0.36492
RV-v5.1 [10] + APL 0.09215 / 0.14391 0.09033 / 0.14496
SDXL [14] 0.27917 / 0.33371 0.27218 / 0.32830
SDXL [14] + APL 0.07267 / 0.14245 0.07852 / 0.14368
Figure 3: Images generated by three different models when employing the same Anonymization Prompts (AP). Models are prompted to generate Albert Einstein and Taylor Swift, which respectively belong to the training set and the testing set. Note that the AP is only trained on SD-v1.5 [1]. The visual effect demonstrate successful anonymization for any identities across different text-to-image models, consistent with our quantitative results.
Following the training of the Anonymization Prompt, we prepend the learned prefix to different text-to-image diffusion models to evaluate its anonymization performances. We prompt the models to generate images of public figures and calculate identity accuracies (ID-ACC). In this experiment, we use the prompt template "A close-up portrait of [Name]" to ensure that the models generate clear depictions of faces. For each identity, we generate a group of images and calculate both mean and maximum ID-ACCs within the group. Subsequently, we aggregate these scores across all identities to present overall mean and maximum identity accuracies. Results are presented in table 1.
Experiments show that original text-to-image models perform well on generating facial identities, reaching an average ID-ACC of around 0.3. However, with the Anonymization Prompts, ID-ACCs drop to as low as 0.039 with SD-v1.5, marking almost a tenfold decrease from the original version. Comparable results of less than 0.1 are obtained on three other models, which is sufficient to render the identities unrecognizable. Remarkably, the anonymization performances on the unseen testing set are on par with those on the training set, indicating an effectiveness to all identities, beyond those available during training.
Visually, faces generated by all models are significantly less recognizable, as demonstrated in  fig. 3, consistent with our quantitative results. We thereby confirm the validity of the Anonymization Prompt on four different models, including the state-of-the-art SDXL [14]. This showcases a remarkable transferability of our method, given the prompt is only trained on SD-v1.5.
@@ 4.3 Quality Preservation
We strive to ensure our anonymization is natural and specific to identifiable facial images. When requested to generate specific individuals, models equipped with the Anonymization Prompts should return realistic facial images with false identities, but retain facial attributes similar to those of the requested identity, as outlined in section 3.4. In addition, this modification should not interfere with generation qualities of unrelated images, i.e., concatenating the Anonymization Prompt to prompts of non-facial images and non-identity-specific facial images should not significantly change the outputs.
@@
In this section, we comprehensively evaluate the ability of our method to meet these requirements. We first study the quality of 4000 facial images generated with identity-specific prompts from our dataset $S_{test}$. We calculate the FID [30] score of our generated faces against those generated by the original model. For attribute evaluation, we calculate the Attribute Accuracy (Attr-ACC) of generated facial images, which is the average cosine similarity between the facial attribute features [33] of the generated faces and the ground truths. For non-identity-specific generations, we generate $30k$ images with prompts from COCO-30k [31]. We evaluate FID scores between images generated by modified models and the original Stable Diffusion [1], and evaluate CLIP scores by averaging the distances between all generated images and their corresponding text prompts.
@@
A direct baseline of our task is to apply concept editing methods [8, 9, 7, 16] and remove facial identities as target concepts. We compare our results with state-of-the-art concept editing, namely ESD [8], Concept Ablation (CA) [9], UCE [16] and Forget-me-not (FMN) [7]. We run their open training code on SD-v1.5 [1] but substitute their editing concepts with identities in our training set.
Table 2: Quantitative results of quality preservation. We report performance on ID-specific generation and non-ID-specific generation, which are respectively evaluated with ID-specific prompts and prompts from COCO-30k [31]. Best results are in bold and second bests are underlined.
ID-specific Non-ID-specific
FID (↓) ID-ACC (↓) Attr-ACC (↑) FID (↓) CLIP (↑)
SD-v1.5 [1] - 0.30254 0.69361 - 0.3078
FMN [7] 263.57 - - 54.89 0.1642
UCE [16] 125.67 - - 88.69 0.0298
ESD [8] 189.32 - - 9.59 0.2530
CA [9] 57.58 0.01215 0.53898 4.14 0.2903
Ours (APL) 20.01 0.03970 0.60222 2.98 0.3058
@@
Results in table 2 demonstrate that our method best satisfies all requirements for a successful anonymization. It not only generates high quality facial images with low identity accuracies, but also preserves facial attribute information of the requested identities. On generating COCO [31] images, APL achieves the highest image quality and text fidelity, respectively quantified by FID and CLIP scores, surpassing all compared methods.
In contrast, the compared methods fail to meet the demand of our anonymization task. After training with our dataset, some methods exhibit difficulties in generating detectable faces, therefore we do not report their ID-ACC and Attr-ACC. For example, FMN [7] and UCE [16] produce high FID scores when prompted with specific identities, such that a large number of images do not contain any faces. Simultaneously, they generate COCO images with high FID scores and low CLIP scores, as the resulting images are neither realistic nor related to the input prompts. This is outcome is consistent with existing research [16], which suggests that editing too many concepts with existing methods may significantly impact the generation quality.
Meanwhile, we show that some methods fail to meet the requirements despite relatively maintaining generation quality. For example, models trained by ESD [8] generates random scenery images when prompted with specific identities, as shown in fig. 4. This occurs because ESD [8] aligns its target concepts to empty strings, effectively removing the concepts entirely. Similarly, Concept Ablation [9] selects anchor concepts, such as "a person", as optimization targets. Although it generates facial images with low identity accuracies, this is at the cost of minimal facial attribute similarity with the requested identity, as indicated by a markedly lower Attr-ACC. As discussed in section 3.4, we consider such anonymization excessive, as they produce completely irrelevant faces or no faces at all. We visualize excessive anonymization in fig. 4.
Figure 4: Excessive anonymization observed in compared methods. Empty string targets of ESD [8] lead to random scenery generations and category label targets of Concept Ablation [9] generate every identity as a random person. Our prompts curated with abundant attribute information strikes a balance between erasure of semantics and attaining original attributes.
@@
APL sets appropriate optimization targets that specifies basic demographic and facial attribute information for all training identities, and therefore preserves image quality and necessary attributes while removing identity information. Its regularization loss $L_{Reg}$ effectively preserves generation quality of unrelated images, as we show in the ablation study. Finally, its prompt learning setting not only ensures minimal quality degradation compared to the original model, but also makes it the only method that can be trained once and then transferred to other models.
@@ 4.4 Anonymizing New Identities
Personalization methods, such as Dreambooth [34], Textual Inversion [35] and HyperDreambooth [5], enable diffusion models to acquire new concepts, including highly accurate facial identities [5]. This capability significantly expands deepfake threats, affecting individuals beyond those in the model’s training dataset. Meanwhile, these methods also undermine the effectiveness of traditional concept editing methods [8, 9], as studies show that concepts removed by them can be re-introduced to the model with the same personalization procedure [6, 36].
Our experiments demonstrate that our method remains effective amid these challenges. Firstly, it effectively anonymizes identities learned with Dreambooth, even though the identities was never seen during training. Secondly, because APL does not modify model parameters, re-learning anonymized identity is not effective. After any attempts of re-learning, users can simply add the prompt back on and still generate anonymized images. These capabilities are not preserved by previously compared methods, which unanimously modify model parameters.
Figure 5: We evaluate the anonymization performance of our Anonymization Prompt (AP) on identites learned after APL training. (a) Identity accuracies of images generated with and without the Anonymization Prompt at different Dreambooth [34] iterations. The prompt consistently lowers identity accuracies on newly-learned identities. (b) An example of anonymizing a new identity [5], which is linked to a token [V] by Dreambooth [34]. The prompt significantly reduces the identity accuracy of the generated image.
In specific, we introduce new identities to Stable Diffusion v1.5 [1] with Dreambooth [34] and evaluate the anonymization performance of the Anonymization Prompt on these identities. We sample 50 identities from the VGGFace-2 [37] dataset, collecting high-resolution photos for each identity for personalization. We obtain a personalized Stable Diffusion for each of those identities with Dreambooth [34], so that resulting models are able to accurately generate their corresponding identities with a pre-determined token. Models are fine-tuned for 800 iterations, which is sufficient for the base model to learn accurate identities. During training, we generate images of the corresponding identities every 100 iteration, both with and without applying the Anonymization Prompt, to measure their identity accuracies. The results across all identities are then aggregated to form the curves in fig. 5(a).
Owing to the effectiveness of Dreambooth [34], both curves exhibit a clear increasing trend. However, in cases where identities are generated with the Anonymization Prompt, the identity accuracies consistently remain lower compared to those generated without the Anonymization Prompt across all iterations. With fewer iterations, the prompt effectively reduces the identity accuracies to nearly half of the base model. At 800 iteratieons, generations with the prompt achieves an average identity accuracy of 0.43, markedly lower than the baseline of 0.53, demonstrating effective anonymizations on these new identities when Dreambooth [34] is trained until convergence. We visualize the effect of the Anonymization Prompt on one learned identity in fig. 5(b). At the same training iteration of Dreambooth [34], the identity generated without the prompt closely resembles the ground truth identity, whereas the one generated with the prompt is much less recognizable, showcasing successful anonymization.
These figures demonstrate both quantitatively and qualitatively that our Anonymization Prompt effectively anonymizes identities learned by the base model after the initial prompt learning. Consequently, modifications to the model parameters to either introduce new identities or restore previously erased ones do not affect our anonymization performance. This finding significantly extends the applicability of APL to encompass all possible facial identities, thereby offering protection against deepfake threats and privacy violations not only for public figures but also for the general population.
@@ 4.5 Ablation Study
We study the effects of several key design choices and hyperparameter selections for Anonymization Prompt Learning, including training iterations, regularization, prompt length, weight parameter $\alpha$ and the size of the training dataset.
@@
Figure 6: Effects of training iterations. (a) An identity generated by RealisticVision [10] that employs the Anonymization Prompts of different training iterations. Different anonymization levels of the same identity is observed. (b) Identity accuracies computed across the test set. Users of our method may choose to use the Anonymization Prompts at different training iterations according to their own policies of content moderation.
@@
Training iterations. We monitor the identity accuracies during the entire training process, and observe that the accuracies for all models progressively decrease over time in fig. 6(a). This effect is visualized with one identity in fig. 6(b). Both visual effects and quantitative measures indicate a steadily increasing level of anonymization. In light of these findings, text-to-image service providers may choose to employ the Anonymization Prompts at various stages of training to enforce different levels of anonymization in alignment with their content moderation policies.
@@
Figure 7: Effect of our regularization loss $L_{Reg}$. When training APL solely with facial images, the model links the prompt to the concept of faces, leading to unexpected depictions of people in every generation. Regularization removes the issue by inactivating the Anonymization Prompt for non-ID prompts.
@@
Regularization. We ablate the regularization loss $L_{Reg}$, which trains the Anonymization Prompt to be ineffective when concatenated to non-identity-specific prompts. We demonstrate in fig. 7 that without regularization, models generate unexpected depictions of people, which is a significant violation of semantic alignment. Since APL is a prompt learning method, training without the regularization of non-face images causes the model to link the prompt to the concept of faces, resulting in above situations. Therefore, training the prompt to ensure its ineffectiveness on non-identity-specific prompts is essential for our method to be useful in practice.
@@
Table 3: Identity Accuracy of the Anonymization Prompt with different prompt length $m$. We report mean / maximum ID-ACC in a group of image for the same individual, and then aggregate across all identities.
@@
Prompt length. The length of the Anonymization Prompt is the number of learnable token that constitutes our prompt prefix. We study the optimal length of the prompt by evaluating Identity Accuracies with prompts of different lengths. As shown in  table 3, we find that $m=10$ strikes a balance between effective anonymization and computational efficiency. Shorter prompts are less effective at anonymization. Longer prompts only offer marginal improvements, while requiring more computational and storage resources.
Table 4: Identity Accuracy of the Anonymization Prompt trained with different weight parameter $\alpha$. We report mean / maximum ID-ACC in a group of image for the same individual, and then aggregate across all identities.
@@
Figure 8: We evaluate the identity accuracies (ID-ACC) of Stable Diffusion v1.5, enhanced with the Anonymization Prompts (AP) trained on differing sizes of datasets. Zero denotes evaluations without the Anonymization Prompt. Evaluation is performed on the test set $S_{test}$ distinct from training identities. We document both the mean and maximum ID-ACC for a group of images per identity, then aggregate these metrics across all identities. We observe a consistent decline in both mean and maximum ID-ACC as the number of identities in the training set increases.
Dataset Size. The Anonymization Prompt enables the anonymization of any identities, including those unseen during training and those learned through fine-tuning. We show that this generalization ability has its roots in the number of training identities. Results are presented in  fig. 8. With more training identities, we observe a consistent decrease in both average and maximum Identity Accuracies. This suggests that incorporating more identities enables the model to identify more precise directions from $c_{1}$ to $c_{2}$ in the latent space, where $c_{1}$ represents the prompts specifiying identities and $c_{2}$ are prompts encapsulating the demographic information and facial attributes of these identities. An accurate estimation of this direction guides the model to anonymize more identities successfully. We expect the identity accuracies to further drop on unseen diffusion models if larger datasets with more diverse attributes are available.
5 Conclusions
We propose Anonymization Prompt Learning (APL), which produces a plug-and-play Anonymization Prompt that can be inserted to any text-to-image diffusion models and enforce the generation of anonymized faces. We show the ability of the Anonymization Prompt to anonymize various identities on different models while preserving the basic generation performances of the original diffusion model. Put together, our method has the potential to be incorporated by text-to-image service platforms to prevent the generation of identity-specific images and address deepfake generations. We hope the development of the method contributes to the long-term battle against malicious use of AI-generated content.
Appendix A Samples from collected dataset
We present some samples from our collected dataset $S_{ID}$ in table 5. We reiterate that the attribute information is automatically generated by ChatGPT [27], and they may not accurately reflect the actual characteristics of these individuals.